# Convergence Rates for Stochastic Proximal and Projection Estimators

### v1

@@ Abstract.
In this paper, we establish explicit, non-asymptotic convergence rates for the stochastic smooth approximations of infimal convolutions introduced and developed in [5, 6]. In particular, we quantify the convergence of the associated barycentric estimators toward proximal mappings and metric projections. We prove a dimension-explicit $\sqrt{\delta}$ bound in the $\rho$-weakly convex (possibly nonsmooth) setting and show, by examples, that this order is sharp. Under additional regularity, namely $C^{2}$ smoothness with globally Lipschitz Hessian, we derive an improved linear $O(\delta)$ rate with explicit constants, and we obtain refined projection estimates for convex sets with local $C^{2,1}$ boundary.
@@ 2. Mathematical Preliminaries
Let $h:\mathbb{R}^{n}\to[0,\infty)$ be an integrable log-concave function, and let
$b_{h}:=\frac{\int_{\mathbb{R}^{n}}y\,h(y)\,dy}{\int_{\mathbb{R}^{n}}h(y)\,dy}$
denotes its barycenter. Then
@@ 3. Nonsmooth results
Claim 1: $\mathbb{E}_{\sigma_{\delta}}\left[\langle y-\operatorname{prox}_{\lambda f}(x),\nabla g(y)\rangle\right]=n\delta$.
Proof of Claim 1: For $R>0$, apply the divergence theorem on $\mathbb{B}_{R}$:
$\int_{\mathbb{B}_{R}}\operatorname{div}F(y)dy=\int_{\partial\mathbb{B}_{R}}\langle F(y),\nu(y)\rangle dS(y).$
Using (2), the right-hand side is bounded in absolutely value by
$CR^{n}\exp\left(-\frac{\mu}{2\delta}(R-\|\operatorname{prox}_{\lambda f}(x)\|)^{2}\right)\to 0\textrm{ as }R\to\infty,$
hence $\int_{\mathbb{R}^{n}}\operatorname{div}F(y)dy=0$. Since $g$ is convex, it is differentiable a.e. and
$\operatorname{div}F(y)dy=ne^{-g(y)/\delta}-\frac{1}{\delta}\langle y-\operatorname{prox}_{\lambda f}(x),\nabla g(y)\rangle e^{-g(y)/\delta}\textrm{ for a.e. }y\in\mathbb{R}^{n}.$
Finally, integrating over $\mathbb{R}^{n}$ and dividing by $\int_{\mathbb{R}^{n}}e^{-g(y)/\delta}dy$, we obtain the claim.
@@ 4. Smooth results
4.1. Stochastic Approximation of Projections onto Smooth Convex Sets
Let $C\subset\mathbb{R}^{n}$ be a nonempty closed convex set with nonempty interior. Recall that
@@ 4.1. Stochastic Approximation of Projections onto Smooth Convex Sets
Let $C\subset\mathbb{R}^{n}$ be a closed set and let $p\in\partial C$. We say that $\partial C$ admits a local $C^{2,1}$ chart at $p$ with constants $(\rho,L,M)$ if, after a rigid motion sending $p$ to 0 and the outward unit normal at $p$ to $e_{n}$, there exists $\rho>0$ and a concave function $h\in C^{2,1}(\mathbb{B}^{n-1}_{\rho})$ such that $h(0)=0$, $\nabla h(0)=0$, $\|\nabla^{2}h(z)\|\leq L$,
@@
Proposition 4.3.
Let $C\subset\mathbb{R}^{n}$ be a closed convex set with nonempty interior, and fix $x\in\mathbb{R}^{n}\setminus C$. Assume that $\partial C$ admits a local $C^{2,1}$ chart at $\operatorname{proj}_{C}(x)$ with constants $(\rho,L,M)$. Then there exists $\delta_{0}>0$ such for every $0<\delta<\delta_{0}$, we have
$\|p_{\delta}(x)-\operatorname{proj}_{C}(x)\|\leq$ $K_{n}\delta,$
where $K_{n}$ is a constant depending on the dimension $n$, $\rho$, $K$, and $M$.
@@
Consider the half-space $C=\{y\in\mathbb{R}^{n}:\langle y,n\rangle\leq 0$, where $\|n\|=1$. Then $\partial C=\{y\in\mathbb{R}^{n}:\langle y,n\rangle=0\}$ is a hyperplane. Pick $x\notin C$, so that $\langle x,n\rangle>0$. Then
$\operatorname{proj}_{C}(x)=x-\langle x,n\rangle n,\quad d_{C}(x)=\langle x,n\rangle.$
@@
$p_{\delta}(x)-\operatorname{proj}_{C}(x)=\alpha_{\delta}\,n$
@@
$p_{\delta}(x)=\operatorname{proj}_{C}(x)+\left(d_{C}(x)-\sqrt{\delta}\frac{\phi(d_{C}(x)/\sqrt{\delta})}{\Phi(-d_{C}(x)/\sqrt{\delta})}\right)n,$

### v2

@@ Abstract.
In this paper, we establish explicit convergence rates for the stochastic smooth approximations of infimal convolutions introduced and developed in [9, 10]. In particular, we quantify the convergence of the associated barycentric estimators toward proximal mappings and metric projections. We prove a dimension-explicit $\sqrt{\delta}$ bound, with explicit constants for the proximal mapping, in the $\rho$-weakly convex (possibly nonsmooth) setting, and we also obtain a dimension-explicit $\sqrt{\delta}$ rate for the metric projection onto an arbitrary convex set with nonempty interior. Under additional regularity, namely $C^{2}$ smoothness with globally Lipschitz Hessian, we derive an improved linear $O(\delta)$ rate with explicit constants, and we obtain refined projection estimates for convex sets with local $C^{2,1}$ boundary. Examples demonstrate that these rates are optimal.
@@ 2. Mathematical Preliminaries
Let $h:\mathbb{R}^{n}\to[0,\infty)$ be a log-concave function, not identically zero, such that $\int_{\mathbb{R}^{n}}h(y)\,dy<\infty$. Define its barycenter by
$b_{h}:=\frac{\int_{\mathbb{R}^{n}}y\,h(y)\,dy}{\int_{\mathbb{R}^{n}}h(y)\,dy}.$
Then $b_{h}$ is well-defined and
@@ 3. Nonsmooth results
Claim 1: $\mathbb{E}_{\sigma_{\delta}}\left[\langle y-\operatorname{prox}_{\lambda f}(x),\nabla g(y)\rangle\right]\leq n\delta$.
Proof of Claim 1: Since $y\mapsto g(y)$ is proper, lsc, and convex, its domain is convex; in particular, $D:=\operatorname{dom}f=\operatorname{dom}g$ is convex with nonempty interior (by assumption). On $\operatorname{int}(D)$, every proper lsc convex function is locally Lipschitz, hence differentiable a.e. and belongs to $W^{1,1}_{\operatorname{loc}}(\operatorname{int}(D))$. Consequently, $\nabla g\in L^{1}_{\operatorname{loc}}(\operatorname{int}(D);\mathbb{R}^{n})$, and for a.e. $y\in\operatorname{int}(D)$,
$\operatorname{div}F(y)=ne^{-g(y)/\delta}-\frac{1}{\delta}\langle y-\operatorname{prox}_{\lambda f}(x),\nabla g(y)\rangle e^{-g(y)/\delta}.$
Fix $R>0$ and set $\Omega_{R}:=\operatorname{int}(D)\cap\mathbb{B}_{R}$. Since
$\|y-\operatorname{prox}_{\lambda f}(x)\|\leq R+\|\operatorname{prox}_{\lambda f}(x)\|=:C_{R}\textrm{ on }\mathbb{B}_{R}$
and $e^{-g(y)/\delta}\leq e^{-\inf_{\mathbb{B}_{R}}g/\delta}=:M_{R,\delta}<\infty$, it follows that for a.e. $y\in\Omega_{R}$,
$|\operatorname{div}F(y)|\leq nM_{R,\delta}+\frac{M_{R,\delta}C_{R}}{\delta}\|\nabla g(y)\|,$
Since $\nabla g\in L^{1}_{\operatorname{loc}}(\operatorname{int}(D);\mathbb{R}^{n})$, we have $\nabla g\in L^{1}(\Omega_{R};\mathbb{R}^{n})$, and then $\operatorname{div}F\in L^{1}(\Omega_{R})$. For $R>0$, apply the divergence theorem on $\Omega_{R}$:
$\int_{\Omega_{R}}\operatorname{div}F(y)dy$ $=\int_{\partial\Omega_{R}}F\cdot\nu\,d\mathcal{H}^{n-1}.$
Moreover, by Lemma A.1, we can select $R_{k}\to\infty$ such that
$\int_{\partial\Omega_{R_{k}}}\langle F,\nu\rangle\,d\mathcal{H}^{n-1}=\underbrace{\int_{\partial D\cap\mathbb{B}_{R_{k}}}F\cdot\nu_{D}\,d\mathcal{H}^{n-1}}_{=:I_{1}}+\underbrace{\int_{\partial\mathbb{B}_{R_{k}}\cap D}F\cdot\nu_{\mathbb{B}_{R_{k}}}\,d\mathcal{H}^{n-1}}_{=:I_{2}}.$
Now, on the one hand, since $D$ is convex and $\operatorname{prox}_{\lambda f}(x)\in D$, for $\mathcal{H}^{n-1}$-a.e. $y\in\partial D$ one has
$\langle y-\operatorname{prox}_{\lambda f}(x),\nu_{D}(y)\rangle\geq 0\quad\Rightarrow\quad\langle F(y),\nu_{D}(y)\rangle\geq 0.$
Hence the $\partial D$-term $I_{1}$ is nonnegative. On the other hand, using (2), the term $I_{1}$ is bounded in absolutely value by
$CR_{k}^{n}\exp\left(-\frac{\mu}{2\delta}(R_{k}-\|\operatorname{prox}_{\lambda f}(x)\|)^{2}\right)\to 0\textrm{ as }k\to\infty,$
hence the $\partial\mathbb{B}_{R_{k}}$-term $I_{2}$ vanishes as $k\to\infty$. Therefore,
$0$ $\leq\limsup_{k\to\infty}\int_{\mathbb{B}_{R_{k}}\cap\partial D}F\cdot\nu_{D}\,d\mathcal{H}^{n-1}$
$=\limsup_{k\to\infty}\Big(\int_{\partial D\cap\mathbb{B}_{R_{k}}}F\cdot\nu_{D}\,d\mathcal{H}^{n-1}+\int_{\partial\mathbb{B}_{R_{k}}\cap D}F\cdot\nu_{\mathbb{B}_{R_{k}}}\,d\mathcal{H}^{n-1}\Big)$
$=\limsup_{k\to\infty}\int_{\partial\Omega_{R_{k}}}F\,\nu\,d\mathcal{H}^{n-1}$
$=\limsup_{k\to\infty}\int_{\Omega_{R_{k}}}\operatorname{div}F(y)dy$
$=\int_{D}\operatorname{div}F(y)dy$
$=n\int_{D}e^{-g(y)/\delta}dy-\frac{1}{\delta}\int_{D}\langle y-\operatorname{prox}_{\lambda f}(x),\nabla g(y)\rangle e^{-g(y)/\delta}dy.$
Finally, dividing by $\int_{D}e^{-g(y)/\delta}dy$ and rearranging, we obtain the claim.∎
@@ 4. Smooth results
4.1. Stochastic Approximation of Projections onto Smooth Sets
Let $C\subset\mathbb{R}^{n}$ be a nonempty closed set with nonempty interior. Recall that
@@ 4.1. Stochastic Approximation of Projections onto Smooth Sets
Let $C\subset\mathbb{R}^{n}$ be a closed convex set and let $p\in\partial C$. We say that $\partial C$ admits a local $C^{2,1}$ chart at $p$ with constants $(\rho,L,M)$ if, after a rigid motion sending $p$ to 0 and the outward unit normal at $p$ to $e_{n}$, there exists $\rho>0$ and a concave function $h\in C^{2,1}(\mathbb{B}^{n-1}_{\rho})$ such that $h(0)=0$, $\nabla h(0)=0$, $\|\nabla^{2}h(z)\|\leq L$,
@@
Theorem 4.3.
Let $C\subset\mathbb{R}^{n}$ be a closed convex set with nonempty interior, and fix $x\in\mathbb{R}^{n}\setminus C$. Assume that $\partial C$ admits a local $C^{2,1}$ chart at $\operatorname{proj}_{C}(x)$ with constants $(\rho,L,M)$. Then there exist $\mathcal{C}>0$ and $\delta_{0}>0$ such for every $0<\delta<\delta_{0}$, we have
$\|p_{\delta}(x)-\operatorname{proj}_{C}(x)\|\leq$ $\,\mathcal{C}\delta+\mathcal{O}(\delta^{2}).$
The next example shows that the order of the bound obtained in Theorem 4.3 is sharp.
@@
Consider the half-space $C=\{y\in\mathbb{R}^{n}:\langle y,n\rangle\leq 0\}$, where $\|\nu\|=1$. Then $\partial C=\{y\in\mathbb{R}^{n}:\langle y,\nu\rangle=0\}$ is a hyperplane. Pick $x\notin C$, so that $\langle x,\nu\rangle>0$. Then
$\operatorname{proj}_{C}(x)=x-\langle x,\nu\rangle\nu,\quad d_{C}(x)=\langle x,\nu\rangle.$
@@
$p_{\delta}(x)-\operatorname{proj}_{C}(x)=\alpha_{\delta}\,\nu$
@@
$p_{\delta}(x)=\operatorname{proj}_{C}(x)+\left(d_{C}(x)-\sqrt{\delta}\frac{\phi(d_{C}(x)/\sqrt{\delta})}{\Phi(-d_{C}(x)/\sqrt{\delta})}\right)\nu,$